OpenStudy (anonymous):

Write the equation of a line with slope m=1/2 and including point (4, -1). y = -2x + 3 y = 2x + 3 y = 1/2x + 3 y = 1/2x - 3

jimthompson5910 (jim_thompson5910):

The slope of y = 2x + 3 is 2, so choice B is out

jimthompson5910 (jim_thompson5910):

same with choice A, the slope is -2. So A is out

OpenStudy (anonymous):

d?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

y = 1/2x - 3 -1 = (1/2)(4) - 3 -1 = 2 - 3 -1 = -1 So the point (4, -1) is on the line y = 1/2x - 3